?1/x^2cos(1/x)dx

Grade:

3 Answers

Badiuddin askIITians.ismu Expert
13 years ago

Dear margaret

1/x2 cos(1/x)dx

put 1/x  =t

      dx/x2 =-dt

intigral become

cos(t)dt

=sint + c

=sin(1/x) + c

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin

margaret miriga
13 years ago

1. Solve the initial value problem:  dy/dx=1/√x+2,  y(2)=-1

 

 

 

sahil malhotra
13 years ago

∫1/x2 cos(1/x)dx put 1/x =t dx/x2 =-dt intigral become ∫ cos(t)dt =sint + c =sin(1/x) + c